# Rigorously Characterizing Dynamics with Machine Learning

### v1

@@ Abstract.
The identification of dynamics from time series data is a problem of general interest. It is well established that dynamics on the level of invariant sets, the primary objects of interest in the classical theory of dynamical systems, is not computable. We recall a coarser characterization of dynamics based on order theory and algebraic topology and prove that this characterization can be identified using approximations that can be realized by feedforward neural networks.
@@ 1. Introduction
The goal of this paper is to present a framework that using machine learning allows for the rigorous characterization of the nonlinear dynamics generated by iterates of a continuous function $f$ defined on a compact subset $X$ of ${\mathbb{R}}^{d}$. To focus the discussion we present three theorems and discuss their relevance.
The first theorem is a classical result from machine learning, the universal approximation theorem [16, 22].
Theorem 1.1.
The premise of this paper relies on the existence of the function $G$ from Theorem 1.1. To simplify the analysis we add the restriction that the activation functions are non-constant and piecewise linear (the results of [22] guarantee that Theorem 1.1 still apply), which allows us to work with approximating functions of the following form.
Theorem 1.2 (Grigsby and Lindsey [12]).
Our strategy is to use $G$ to characterize the dynamics of $f$. The classical theory of dynamical systems focuses on understanding the existence and structure of invariant sets, i.e., sets of the form $S\subset X$ such that $f(S)=S$. This leads to a rich mathematical theory but, as we argue below, a theory that is too rich from the perspective of deducing dynamics from data.
With regard to the goals of this paper, the fundamental question is: is there a $\rho$ such that $G$ (satisfying the conditions of Theorem 1.1) generates the same dynamics as $f$? In the classical setting this is equivalent to asking: can we guarantee that $f$ and $G$ are conjugate, i.e., does there exist a homeomorphism $h\colon X\to X$ such that
@@
The work of [9] proves that in general this is not possible via computations. More precisely, regarding their work the authors state
@@
A trivial example that demonstrates the challenge of identifying invariant sets based on approximations is as follows.
Consider $f_{\theta}\colon{\mathbb{R}}\to{\mathbb{R}}$ defined by
@@
Figure 1. (a) The function $f_{\theta}$ for $\theta<1$ that has a fixed point at 00. (b) The function $f_{1}$ that has fixed points at 00 and $1$. (c) The function $f_{\theta}$ for $\theta>1$ that has fixed points at 00, $1$, and $\theta$.
Returning to the goal of this paper, we are allowed to assume that the map we want to learn the dynamics of is $f_{1}\colon{\mathbb{R}}\to{\mathbb{R}}$, but we are not allowed to assume that we know the functional form of $f_{1}$. We are also allowed to assume that we have a finite (but arbitrarily large) set of data $\left\{{(x_{k},f_{1}(x_{k}))\mid x_{k}\in[-2,2]}\right\}$ that can be used to train $G\colon{\mathbb{R}}\to{\mathbb{R}}$, and that as a result of this training, $\sup_{x\in[-2,2]}|f_{1}(x)-G(x)|<\rho$. Observe that even for arbitrarily small $\rho>0$, we cannot guarantee that $G$ has exactly two fixed points, and therefore we cannot guarantee that we can learn the number of fixed points of $f_{1}$ from $G$.
While Example 1.3 may appear to be too trivial to be of interest, it is indicative of a more serious challenge. Consider a continuous parameterized family of maps $f\colon X\times[0,1]^{n}\to X$, where for each $\lambda\in[0,1]^{n}$, $f_{\lambda}:=f(\cdot,\lambda)$. A parameter value $\lambda_{0}$ is a bifurcation point if for any neighborhood $U$ of $\lambda_{0}$ there exists $\lambda\in U$ such that $f_{\lambda}$ is not conjugate to $f_{\lambda_{0}}$. In particular, $\theta=1$ is a bifurcation point in Example 1.3. As is shown in [33, 34] if $d\geq 2$ (recall that $X\subset{\mathbb{R}}^{d}$), then there are parameterized families of maps for which the bifurcation points contain a Cantor set of positive measure. As a consequence, in general it is not possible to identify the dynamics of $f$ via an approximation provided by Theorem 1.1.
We avoid the above mentioned challenges by employing a coarser framework for dynamics that we believe to be sufficiently refined for many applications. Since we do not expect that the typical reader is familiar with this framework and its associated machinery, a significant portion of this paper is spent introducing the necessary definitions and concepts. In particular, in Section 2 we provide precise definitions of the terms used in the following theorem, which is our main result.
@@
Let $G\colon{\mathbb{R}}^{d}\to{\mathbb{R}}^{d}$ be a fully-connected feedforward neural network with piecewise linear activation such that $\sup_{x\in X}\|f(x)-G(x)\|<\rho$. If $\rho>0$ is sufficiently small, then using $G$ we can
@@
An imprecise statement of Theorem 1.4 is as follows. Choose a Morse representation ${\mathsf{MR}}$ (see Definition 2.4) for the function $f$. In general $f$ has many (possibly countably infinitely many) Morse representations, thus the choice of ${\mathsf{MR}}$ can be interpreted as a choice of resolution of the dynamics of $f$ that is of interest. The theorem implies that we can employ a neural network $G$ to (i) Organize the global dynamics of $f$ in terms of its gradient-like behavior, and (ii) Compute homological invariants (Conley indices) from which we can obtain rigorous statements concerning the dynamics of $f$.
To provide slightly more detail, a Morse graph (see Definition 3.2) is a partially ordered set (poset) whose elements identify potential recurrent dynamics and whose order structure indicates the asymptotic direction of nonrecurrent dynamics. The Conley index ${\mathsf{Con}}_{*}$ (see Definition 2.9) is a computable [31] algebraic topological extension of the Morse index that we employ to identify nonlinear dynamics. Most fundamentally, a non-trivial Conley index implies the existence of a non-empty invariant set, however the converse is not true. The Conley index can also be used to prove the existence of equilibria [39, 24], periodic orbits [27], heteroclinic orbits [4], chaotic dynamics [29, 42, 8, 7], and semi-conjugacies onto nontrivial dynamics [28, 26, 25]. Thus, the framework being considered in this paper is capable of identifying many of the building blocks of traditional nonlinear dynamics.
@@
Theorem 1.4 guarantees that given the constraint that $\rho$ is sufficiently small there is a constructive method for identifying the desired dynamics. This may appear to be a weak result since no quantitative bounds are provided for $\rho$. However, the complexity of the geometry of attractors and the boundaries of their domains of attraction can be extreme [21], and this geometry plays an intrinsic role in the necessary value of $\rho$. Given the level of generality at which we are working (the only essential assumption is that $f$ is continuous) we see no path forward, based on finite data, to provide a sharp quantification of potential values of $\rho$.
@@
We use the family of maps introduced in Example 1.3 to provide intuition about the content of Theorem 1.4. We leave it to the reader to check that $f_{1}([-2,2])=[0,1]\subset\mathop{\mathrm{int}}\nolimits([-2,2])$. Consider (using the language of Theorem 1.4) a Morse representation $({\mathsf{MR}},\leq_{{\mathsf{MR}}_{f_{1}}})$ for $f_{1}$ whose elements, called Morse sets, are the fixed points (invariant sets) $x=0$ and $x=1$, and for which the partial order is $10\leq_{{\mathsf{MR}}_{f_{1}}}$ 1. The partial order captures the fact that if $x\in(0,1)$ then $\lim_{n\to\infty}f_{1}^{n}(x)=0$ and $\lim_{n\to-\infty}f_{1}^{n}(x)=1$. The Morse sets identify the recurrent dynamics of $f_{1}$ and the partial order characterizes the non-recurrent (gradient-like) dynamics of $f_{1}$. The Conley index of 00 is that of an attracting fixed point, while the Conley index of $1$ is trivial.
@@
As discussed above, in general, invariant sets (Morse sets) are not computable, and therefore, Morse representations are not computable. What are computable are Morse graphs (see Definition 3.2). Theorem 1.4(i) provides the link between the computable Morse graph and the desired Morse representation. The computation of the Morse graph ${\mathsf{MG}}$ is based on Theorem 1.2 that provides a finite polyhedral decomposition of phase space and the ability to easily bound the images of these polytopes under $G$. For this example the polytopes are intervals. This allows us to produce a combinatorial representation of the dynamics in the form of a directed graph ${\mathcal{G}}$ (see Sections 3 and 4) where the vertices correspond to the above mentioned intervals and the directed edges indicate how intervals are mapped amongst each other. The nodes of the associated Morse graph capture recurrence within the directed graph and the order relation provides information about the existence of nonrecurrent paths.
@@
So, assume that $G=f_{\theta}$ for $\theta<1$. By assumption $\sup_{x\in[-2,2]}\|f(x)-G(x)\|<\rho$. Let $I_{0}$ denote the interval that contains 00. Then $G(I_{0})\cap I_{0}\neq\emptyset$. Thus, under $G$ the interval $I_{0}$ that contains 00 clearly maps to itself, and hence the directed graph ${\mathcal{G}}$ contains the self edge $I_{0}\to I_{0}$. Let $I_{1}$ denote the interval that contains $1$. Similarly, (since $\rho$ is assumed small or equivalently $\theta\approx 1$) the above mentioned bounds implies that $G(I_{1})\cap I_{1}\neq\emptyset$ and hence ${\mathcal{G}}$ contains the self edge $I_{1}\to I_{1}$. Ideally, intervals between $I_{0}$ and $I_{1}$ are mapped strictly to the left, in which case the Morse graph ${\mathsf{MG}}$ has the form $I_{1}\rightarrow I_{0}$. The poset morphism $\nu\colon{\mathsf{MG}}\to{\mathsf{MR}}$ has the form $\nu(I_{0})=0$ and $\nu(I_{1})=1$, and in this idealized case $\nu$ is an isomorphism.
As described in Section 4.2, ${\mathcal{G}}$ can be used to compute Conley indices and furthermore these Conley indices are valid for the continuous dynamics generated by $G$. The Conley index associated with $I_{0}$ is that of an attracting fixed point. The Conley index associated with $I_{1}$ is trivial; it must be since under $G$ the invariant set associated with $I_{1}$ is the empty set. The fact that $\nu$ is a poset morphism allows us to conclude that the Conley indices for the fixed points 00 and $1$ of $f_{1}$ are that of an attracting fixed point and trivial, respectively. Referring again to [39, 24], knowledge that the Conley index of $\nu(I_{0})$ is that of a fixed point allows us to conclude that the invariant set associated with $\nu(I_{0})$ contains a fixed point. As emphasized above the fact that the Conley index of $\nu(I_{1})$ is trivial does not imply that the invariant set associated with $\nu(I_{1})$ is empty. The Morse graph ${\mathsf{MG}}$ identified $I_{1}$ and the claim that $\nu$ is an epimorphism implies that $\nu(I_{1})$ is a non-empty invariant set, but suggests that it may not be robust with respect to perturbations.
For a variety of reasons $\nu$ need not be an isomorphism (these reasons are what make the proof of Theorem 1.4 challenging). Again, with the goal of gaining intuition, assume that $G=f_{\theta}$ for $\theta>1$. The explicit form of $G$ (see Figure 1(c)) and our assumption that $G$ satisfies Theorem 1.2 allows us to deduce that there are distinct intervals $I_{0}$, containing 00, $I_{1}$, containing $1$, and $I_{\theta}$, containing $\theta$, on which $G$ is affine-linear. The combinatorial representation of the dynamics of $G$ must contain the self edges $I_{0}\to I_{0}$, $I_{1}\to I_{1}$, and $I_{\theta}\to I_{\theta}$.
@@
As indicated above, ${\mathcal{G}}$ can be used to compute Conley indices. For this case the Conley indices of $I_{0}$ and $I_{\theta}$ are that of a stable fixed point, while the Conley index of $I_{1}$ is that of a fixed point with an unstable manifold of dimension one. However, ${\mathcal{G}}$ also allows us to determine that under $f_{1}$ the Conley indices of 00 and $1$ are those of a stable fixed point and trivial, respectively.
@@
We remind the reader that Theorem 1.4 demonstrates that using the coarser framework of dynamics of this paper, i.e. a Morse representation, dynamics generated by a continuous function can be recovered by a fully-connected feedforward neural network $G$. In the context of applications, we do not know $f$ and hence cannot know a Morse representation $({\mathsf{MR}},\leq_{{\mathsf{MR}}_{f}})$. In practice, $G$ is learned from data, ${\mathcal{G}}$ is constructed, and then conclusions are drawn about the dynamics of $f$. We return to this point in Section 6.
An outline of the paper is as follows. In Section 2, we introduce Conley theory in the context of continuous maps and describe how dynamics can be characterized in this framework. To demonstrate the generalizability of this framework, we prove Theorem 2.25, which loosely states that for any continuous function $g$ sufficiently close to a continuous function $f$, the characterization of dynamics of $f$ also characterizes the dynamics of $g$. Section 3 contains an introduction to combinatorial dynamics and culminates by describing how a Morse graph is computed. In Section 4, we connect this perspective back to continuous dynamics and introduce the piecewise linear approximations that are the focus of this work. In Section 5, we prove the main result of the paper, Theorem 1.4. Section 6 contains concluding remarks about future work and how the ideas of this paper may be used in practice.

### v2

@@ Abstract.
The identification of dynamics from time series data is a problem of general interest. It is well established that dynamics on the level of invariant sets, the primary objects of interest in the classical theory of dynamical systems, is not computable. We recall a coarser characterization of dynamics based on order theory and algebraic topology and prove that this characterization can be identified using approximations.
@@ 1. Introduction
The goal of this paper is to present a framework that allows for the use of machine learning to rigorously characterize nonlinear dynamics generated by iterates of a continuous function $f$ defined on a compact subset $X$ of ${\mathbb{R}}^{d}$. The fundamental assumption is that given sufficient data and given $\rho>0$ it is possible to learn a function $G\colon{\mathbb{R}}^{d}\to{\mathbb{R}}^{d}$ that satisfies the following three properties.
P1:
$\sup_{x\in X}\|f(x)-G(x)\|<\rho$.
P2:
Given $x\in X$ we can evaluate $G(x)$.
P3:
We can determine explicit Lipschitz bounds on $G$.
Our strategy is to use $G$ to characterize the dynamics of $f$.
The classical theory of dynamical systems focuses on understanding the existence and structure of invariant sets, i.e., sets of the form $S\subset X$ such that $f(S)=S$. This leads to a rich mathematical theory but, as we argue below, a theory that is too rich from the perspective of deducing dynamics from data.
With regard to the goals of this paper, the fundamental question is: is there a $\rho$ such that $G$ (satisfying P1 - P3) generates the same dynamics as $f$? In the classical setting this is equivalent to asking: can we guarantee that $f$ and $G$ are conjugate, i.e., does there exist a homeomorphism $h\colon X\to X$ such that
@@
The work of [foreman:rudolph:weiss] proves that in general this is not possible via computations. More precisely, regarding their work the authors state
@@
A trivial example that demonstrates the challenge of identifying invariant sets based on approximations is as follows. Consider $f_{\theta}\colon{\mathbb{R}}\to{\mathbb{R}}$ defined by
@@
Figure 1. (a) The function $f_{\theta}$ for $\theta<1$ that has a fixed point at 0. (b) The function $f_{1}$ that has fixed points at 0 and $1$. (c) The function $f_{\theta}$ for $\theta>1$ that has fixed points at 0, 11, and $\theta$.
Returning to the goal of this paper, we are allowed to assume that the map we want to learn the dynamics of is $f_{1}\colon{\mathbb{R}}\to{\mathbb{R}}$, but we are not allowed to assume that we know the functional form of $f_{1}$. We are also allowed to assume that we have a finite (but arbitrarily large) set of data $\left\{{(x_{k},f_{1}(x_{k}))\mid x_{k}\in[-2,2]}\right\}$ that can be used to obtain an approximation $G\colon{\mathbb{R}}\to{\mathbb{R}}$ such that $\sup_{x\in[-2,2]}|f_{1}(x)-G(x)|<\rho$. Observe that even for arbitrarily small $\rho>0$, we cannot guarantee that $G$ has exactly two fixed points, and therefore we cannot guarantee that we can learn the number of fixed points of $f_{1}$ from $G$.
While Example 1.1 may appear to be too trivial to be of interest, it is indicative of a more serious challenge. Consider a continuous parameterized family of maps $f\colon X\times[0,1]^{n}\to X$, where for each $\lambda\in[0,1]^{n}$, $f_{\lambda}:=f(\cdot,\lambda)$. A parameter value $\lambda_{0}$ is a bifurcation point if for any neighborhood $U$ of $\lambda_{0}$ there exists $\lambda\in U$ such that $f_{\lambda}$ is not conjugate to $f_{\lambda_{0}}$. In particular, $\theta=1$ is a bifurcation point in Example 1.1. As is shown in [newhouse, palis:takens] if $d\geq 2$ (recall that $X\subset{\mathbb{R}}^{d}$), then there are parameterized families of maps for which the bifurcation points contain a Cantor set of positive measure. As a consequence, in general it is not possible to identify the dynamics of $f$ via an approximation.
We avoid the above mentioned challenges by employing a coarser framework for dynamics that we believe to be sufficiently refined for many applications. Figure 2 presents an overview of our approach.
Figure 2. Overview of the proposed framework. (Blue) Given a dynamical system $f$, a Morse representation $\mathsf{MR}$} is equivalent to a choice of resolution of the dynamics. In applications $f$ and $\mathsf{MR}$} are unknown. (Purple) Theory guarantees the existence of a lattice of attracting blocks $\mathsf{N}$} that in turn determines an error bound $\rho$. In applications $\mathsf{N}$} and $\rho$ are unknown. (Pink) Given $\rho$, we assume the existence of an approximation $G$ that satisfies P1. In applications, the approximation $G$ is derived from data generated by $f$. (Orange) Given $G$, we compute a combinatorial model $\mathcal{G}$} that takes the form of a directed graph on a discretization of the phase space $\mathcal{X}$}. Graph algorithms produce a Morse graph, which captures the gradient-like structure of the dynamics. Homology computations produce Conley indices, topological invariants that provide further information about the recurrent dynamics. (Dashed line) Theorem 1.2 guarantees that an order-preserving surjection $\nu$ from the Morse graph to the chosen Morse representation exists. As a consequence the computations provide correct information for any system satisfying the error bound.
Since we do not expect that the typical reader is familiar with this framework and its associated machinery, a significant portion of this paper is spent introducing the necessary definitions and concepts. In particular, in Section 2 we provide precise definitions of the terms used in the following theorem, which is our main result.
@@
Let $G\colon{\mathbb{R}}^{d}\to{\mathbb{R}}^{d}$ satisfy P1 - P3. If $\rho>0$ is sufficiently small, then using $G$ we can
@@
An imprecise statement of Theorem 1.2 is as follows. Choose a Morse representation $\mathsf{MR}$} (see Definition 2.4) for the function $f$. In general $f$ has many (possibly infinitely many) Morse representations, thus the choice of $\mathsf{MR}$} can be interpreted as a choice of resolution of the dynamics of $f$ that is of interest. The theorem implies that we can employ an approximation $G$ to (i) organize the global dynamics of $f$ in terms of its gradient-like behavior, and (ii) compute homological invariants (Conley indices) from which we can obtain rigorous statements concerning the dynamics of $f$.
To provide slightly more detail, a Morse graph (see Definition 3.2) is a partially ordered set (poset) whose elements identify potential recurrent dynamics and whose order structure indicates the asymptotic direction of nonrecurrent dynamics. The Conley index $\mathsf{Con}$$_{*}$ (see Definition 2.9) is a computable [mischaikow:weibel] algebraic topological extension of the Morse index that we employ to identify nonlinear dynamics. Most fundamentally, a non-trivial Conley index implies the existence of a non-empty invariant set, however the converse is not true. The Conley index can also be used to prove the existence of equilibria [srzednicki:85, mccord:88], periodic orbits [mccord:mischaikow:mrozek], heteroclinic orbits [conley:cbms], chaotic dynamics [mischaikow:mrozek:95, szymczak:96, day:junge:mischaikow, day:frongillo], and semi-conjugacies onto nontrivial dynamics [mischaikow:95, McCord:Mischaikow:96, mccord:00]. Thus, the framework being considered in this paper is capable of identifying many of the building blocks of traditional nonlinear dynamics.
@@
Theorem 1.2 guarantees that given the constraint that $\rho$ is sufficiently small there is a constructive method for identifying the desired dynamics. This may appear to be a weak result since no quantitative bounds are provided for $\rho$. However, the complexity of the geometry of attractors and the boundaries of their domains of attraction can be extreme [kennedy:yorke], and this geometry plays an intrinsic role in the necessary value of $\rho$. Given the level of generality at which we are working (the only essential assumption is that $f$ is continuous) we see no path forward, based on finite data, to provide a sharp quantification of potential values of $\rho$.
As presented in Theorem 1.2 our result can be viewed as an abstract approximation theorem. To emphasize that P1 - P3 can (at least conceptually) be realized via some form of machine learning we recall two theorems associated with feedforward neural networks, also known as multilayer perceptrons. The first theorem is a classical result from machine learning, the universal approximation theorem [cybenko, hornik, leshno:lin:pinkus:schocken], which guarantees that $\bf P1$} can be achieved.
Theorem 1.4.
To guarantee that P3 can be realized we add the restriction that the activation functions are non-constant and piecewise linear (the results of [leshno:lin:pinkus:schocken] guarantee that Theorem 1.4 still apply). This gives rise to approximating functions of the following form.
Theorem 1.5 (Grigsby and Lindsey [grigsby:lindsey]).
Observe that theoretically, local Lipschitz constants can be obtained by evaluating $G$ on the cells of $\mathcal{P}$}. The work of [fazlyab2019] provides an algorithm to compute an accurate upper bound for a global Lipschitz constant for $G$.
Remark 1.6.
As suggested in Remark 1.3 in this paper we do not address the question of how to achieve $G$ using a machine learning method. Rather, our focus is on demonstrating that given $G$, it is possible to rigorously identify dynamics of $f$ using the framework that we are proposing.
@@
We use the family of maps introduced in Example 1.1 to provide intuition about the content of Theorem 1.2. We leave it to the reader to check that $f_{1}([-2,2])=[0,1]\subset\mathop{\mathrm{int}}\nolimits([-2,2])$. Consider (using the language of Theorem 1.2) a Morse representation $({\mathsf{MR}},\leq_{{\mathsf{MR}}_{f_{1}}})$ for $f_{1}$ whose elements, called Morse sets, are the fixed points (invariant sets) $x=0$ and $x=1$, and for which the partial order is $0\leq_{{\mathsf{MR}}_{f_{1}}}1$. The partial order captures the fact that if $x\in(0,1)$ then $\lim_{n\to\infty}f_{1}^{n}(x)=0$ and $\lim_{n\to-\infty}f_{1}^{n}(x)=1$. The Morse sets identify the recurrent dynamics of $f_{1}$ and the partial order characterizes the non-recurrent (gradient-like) dynamics of $f_{1}$. The Conley index of 0 is that of an attracting fixed point, while the Conley index of $1$ is trivial.
@@
As discussed above, in general, invariant sets (Morse sets) are not computable, and therefore, Morse representations are not computable. What are computable are Morse graphs (see Definition 3.2). Theorem 1.2(i) provides the link between the computable Morse graph and the desired Morse representation. The computation of the Morse graph $\mathsf{MG}$} is based on the assumption that $X$ is a polygonal region, P2 and P3, and a discretization of the phase space. For this example the discretization is composed of intervals. This allows us to produce a combinatorial representation of the dynamics in the form of a directed graph $\mathcal{G}$} (see Sections 3 and 4) where the vertices correspond to the above mentioned intervals and the directed edges indicate how intervals are mapped amongst each other. The nodes of the associated Morse graph capture recurrence within the directed graph and the order relation provides information about the existence of nonrecurrent paths.
@@
So, assume that $G=f_{\theta}$ for $\theta<1$. By assumption $\sup_{x\in[-2,2]}\|f(x)-G(x)\|<\rho$. Let $I_{0}$ denote the interval that contains 0. Then $G(I_{0})\cap I_{0}\neq\emptyset$. Thus, under $G$ the interval $I_{0}$ that contains 0 clearly maps to itself, and hence the directed graph $\mathcal{G}$} contains the self edge $I_{0}\to I_{0}$. Let $I_{1}$ denote the interval that contains $1$. Similarly, (since $\rho$ is assumed small or equivalently $\theta\approx 1$) the above mentioned bounds implies that $G(I_{1})\cap I_{1}\neq\emptyset$ and hence $\mathcal{G}$} contains the self edge $I_{1}\to I_{1}$. Ideally, intervals between $I_{0}$ and $I_{1}$ are mapped strictly to the left, in which case the Morse graph $\mathsf{MG}$} has the form $I_{1}\rightarrow I_{0}$. The poset morphism $\nu\colon{\mathsf{MG}}\to{\mathsf{MR}}$ has the form $\nu(I_{0})=0$ and $\nu(I_{1})=1$, and in this idealized case $\nu$ is an isomorphism.
As described in Section 4.2, $\mathcal{G}$} can be used to compute Conley indices and furthermore these Conley indices are valid for the continuous dynamics generated by $G$. The Conley index associated with $I_{0}$ is that of an attracting fixed point. The Conley index associated with $I_{1}$ is trivial; it must be since under $G$ the invariant set associated with $I_{1}$ is the empty set. The fact that $\nu$ is a poset morphism allows us to conclude that the Conley indices for the fixed points 0 and $1$ of $f_{1}$ are that of an attracting fixed point and trivial, respectively. Referring again to [srzednicki:85, mccord:88], knowledge that the Conley index of $\nu(I_{0})$ is that of a fixed point allows us to conclude that the invariant set associated with $\nu(I_{0})$ contains a fixed point. As emphasized above the fact that the Conley index of $\nu(I_{1})$ is trivial does not imply that the invariant set associated with $\nu(I_{1})$ is empty. The Morse graph $\mathsf{MG}$} identified $I_{1}$ and the claim that $\nu$ is an epimorphism implies that $\nu(I_{1})$ is a non-empty invariant set, but suggests that it may not be robust with respect to perturbations.
For a variety of reasons $\nu$ need not be an isomorphism (these reasons are what make the proof of Theorem 1.2 challenging). Again, with the goal of gaining intuition, assume that $G=f_{\theta}$ for $\theta>1$. The explicit form of $G$ (see Figure 1(c)) and our assumption that $G$ satisfies P1 - P3 allows us to deduce that there are distinct intervals $I_{0}$, containing 0, $I_{1}$, containing $1$, and $I_{\theta}$, containing $\theta$, on which $G$ is affine-linear. The combinatorial representation of the dynamics of $G$ must contain the self edges $I_{0}\to I_{0}$, $I_{1}\to I_{1}$, and $I_{\theta}\to I_{\theta}$.
@@
As indicated above, $\mathcal{G}$} can be used to compute Conley indices. For this case the Conley indices of $I_{0}$ and $I_{\theta}$ are that of a stable fixed point, while the Conley index of $I_{1}$ is that of a fixed point with an unstable manifold of dimension one. However, $\mathcal{G}$} also allows us to determine that under $f_{1}$ the Conley indices of 0 and $1$ are those of a stable fixed point and trivial, respectively.
@@
We remind the reader that Theorem 1.2 demonstrates that using the coarser framework of dynamics of this paper, i.e., a Morse representation, dynamics generated by a continuous function can be recovered by a fully-connected feedforward neural network $G$. In the context of applications, we do not know $f$ and hence cannot know a Morse representation $({\mathsf{MR}},\leq_{{\mathsf{MR}}_{f}})$. In practice, $G$ is learned from data, $\mathcal{G}$} is constructed, and then conclusions are drawn about the dynamics of $f$. We return to this point in Section 6.
An outline of the paper is as follows. In Section 2, we introduce Conley theory for continuous maps and describe how dynamics can be characterized in this framework. To demonstrate the generalizability of this framework, we present Theorem 2.25, which loosely states that for any continuous function $g$ sufficiently close to a continuous function $f$, the characterization of dynamics of $f$ also characterizes the dynamics of $g$. In Section 3, we introduce combinatorial dynamics and describe how a Morse graph is computed. In Section 4, we connect this perspective back to continuous dynamics and discuss combinatorial representations of dynamics obtained from approximations. In Section 5, we prove Theorem 1.2. We conclude in Section 6 with a demonstration of the applicability of this approach.